# arXiv:2603.01466v1[math.FA] 02 Mar 2026

### v1

@@ Abstract.
Differently from the non-relativistic quantum mechanics, the violation of Bell inequalities in quantum field theory depends more on the structure of observable algebras (typically type III von Neumann algebras) rather than the choice of specific quantum states. Therefore, studying the violation of Bell inequalities based on the von Neumann algebraic framework often reveals information about the algebraic structure. In this paper, we employ three mutually-commuting von Neumann algebras to characterize quantum entanglement swapping networks, and establish Bell-like inequalities thereon, commonly referred to as bilocal inequalities. We investigate the algebraic structural conditions under which bilocal inequalities are satisfied or violated on the generated algebra of these three von Neumann algebras. Furthermore, the conditions for maximal violation of the inequalities can be utilized to infer the structural information of von Neumann algebras in reverse. Our results not only utilize the violation of bilocal inequalities to reveal the structural properties of von Neumann algebras, but can also be applied to quantum mechanics and quantum field theory.
@@ 1. Introduction
Bell nonlocality demonstrates that local measurements performed on one subsystem of a quantum state can instantaneously influence the measurement outcomes on another subsystem, regardless of the spatial separation between them [1, 2, 3]. Such nonlocal correlations can be detected through Bell inequalities, which serve as constraints that all local correlations must obey [4, 5, 6, 7, 8]. It has been demonstrated to offer quantum advantages in various device-independent quantum information tasks, including communication complexity [9], quantum key distribution [10, 11], randomness amplification [12, 13], and measurement-based quantum computation [14, 15].
Moreover, Bell nonlocality in quantum field theory (QFT) serves as a crucial bridge connecting quantum information science with fundamental physics [16, 17]. It provides a rigorous framework for reconciling quantum entanglement with relativistic causality, resolves conceptual challenges such as impossible measurements, and reveals how fundamental symmetries like parity violation affect quantum correlations [19, 18]. Furthermore, the tension between nonlocal quantum phenomena and the locality assumptions of general relativity offers profound insights for quantum gravity research, suggesting that spacetime itself may be an emergent property rather than a fundamental aspect of reality [20, 21, 22].
In QFT, the algebras assigned to touching spacetime regions are of type III von Neumann ones, so there are no pure normal states and no tensor-product structures [23, 24, 25]. This differs from the traditional quantum-mechanical setup, which is usually linked to type I von Neumann algebras and relies on the algebraic tensor product as its mathematical framework [27, 26, 28]. These are two distinct models, referred to respectively as the tensor product algebra model and the mutually-commuting von Neumann algebra model. In the tensor-product algebra model, a bipartite composite quantum system is described by the tensor product of two Hilbert spaces $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$, denoted as $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}.$ Let $\mathcal{B}(\mathcal{H})$ denote the algebra of all bounded linear operators on $\mathcal{H}$, and $\mathcal{T}(\mathcal{H})$ the space of trace-class operators. Then there is a duality relation $\mathcal{B}(\mathcal{H})=\mathcal{T}(\mathcal{H})^{*}$. A quantum state $\rho$ is a positive operator with trace 1, and the set of states is denoted by $\mathcal{S}(\mathcal{H})=\{\rho\in\mathcal{T}(\mathcal{H})\mid\rho\geq 0,\ \mathrm{Tr}(\rho)=1\}$. The algebra of observables is the Jordan algebra of self-adjoint operators, denoted by $\mathcal{B}_{s}(\mathcal{H})=\{A\in\mathcal{B}(\mathcal{H})\mid A=A^{*}\}.$ However, it should be pointed out that relying solely on the above tensor-product model to discuss quantum information problems has two drawbacks [29, 30, 31]: 1) It fails to provide a universal framework for accurately describing phenomena in infinite-dimensional systems. 2) It is not suitable for studying traditional quantum physics such as quantum field theory, which requires the language of type III von Neumann algebras. In the mutually-commuting von Neumann algebra model, the algebra of observables of a quantum system is described by a von Neumann algebra $\mathcal{M}$, where $\mathcal{M}\subseteq\mathcal{B}(\mathcal{H})$, and $\mathcal{M}_{A},\mathcal{M}_{B}$ are two mutually commuting von Neumann subalgebras of $\mathcal{M}$. The notation $\mathcal{M}^{\prime\prime}$ denotes the double commutant of $\mathcal{M}$. However, in more general cases, a natural question arises: are the two models isomorphic? In other words, does $\mathcal{M}$ admit an isomorphism to $\mathcal{M}_{A}\otimes\mathcal{M}_{B}$? This question is intimately related to the famous Tsirelson problem, which is equivalent to the Connes embedding conjecture in von Neumann algebra theory [33, 32, 34]. The Tsirelson problem asks whether the sets of joint probability distributions of measurement outcomes are the same in the two models. In 2020, Ji and collaborators [34] designed a quantum interactive protocol based on the classical halting problem in computability theory, and proved that the answer to the Tsirelson problem is negative. This means that the tensor-product model algebra model and the mutually-commuting von Neumann algebra model are not equivalent. Therefore, the mutually-commuting von Neumann algebra model provides a more general framework, and many novel quantum phenomena fundamentally different from those in the tensor-product model have been discovered in this setting [34, 35, 36, 37].
@@
Overall, Bell nonlocality is not merely a quantum peculiarity but a structural feature encoded in the classification of operator algebras, providing rigorous tools to quantify non-classical correlations in relativistic quantum systems [17]. Summers et al. introduces the maximal Bell correlation and proves that its value is bounded by $\sqrt{2}$ in mutually-commuting von Neumann algebra model, with equality attainable iff each subalgebra contains a copy of $\mathcal{M}_{2}({\mathbb{C}})$ [38]. Translating these bounds into the vacuum representation of algebraic quantum field theory, he shows that tangent wedge algebras are always maximally correlated, whereas strictly spacelike-separated wedges and decay exponentially with mass-governed distance [39, 40, 41, 42, 43]. These works expose a new algebraic invariant¡ªthe Bell correlation invariant¡ªthat distinguishes infinitely many isomorphism classes of mutually-commuting von Neumann algebra pairs and links maximal violation to the presence of the hyperfinite type-${\rm II}_{1}$ factor [17].
@@
In contrast to entanglement originating from an individual source, quantum networks comprise numerous small-scale entangled states. Owing to the independence among distinct sources, the correlations emerging from quantum networks exhibit non-convex characteristics that transcend the polytopes associated with single-source entanglement [44, 45, 46, 47, 48, 49]. To date, Bell-type inequalities have been devised to certify nonlocal correlations across diverse network architectures, such as entanglement-swapping networks [44, 50], chain configurations [51, 52], star topologies [53, 54], polygon structures [55, 56, 57], tree-shaped networks [58, 59, 60], arbitrary acyclic networks [46, 47, 61], and arbitrary $k$-independent networks [62]. Alternative research directions examine the stronger forms of network nonlocality that surpass hybrid implementations involving classical variables and post-quantum resources [63, 64]. Furthermore, the interplay between graph theory and quantum correlations has attracted considerable attention [65, 66]. Nevertheless, limited progress has been made concerning the discrimination of correlations produced by different networks and the subsequent identification of underlying quantum topologies [67]. Interestingly, the notion of bi-locality in the mutually-commuting von Neumann algebra model has already been introduced by Ligthart et al [68], and Xu has addressed the inclusion problem between the tensor product algebra model and the mutually-commuting von Neumann algebra model in this setting [69].
In this paper we aim to establish bi-local inequalities within the mutually-commuting von Neumann algebra model and investigate how the degree of their violation is related to the structural properties of the algebras.
@@ 2. Ternary mutually-commuting von Neumann algebra models and entanglement swapping networks
Quantum bilocal scenario. In non-relativistic quantum mechanics, the quantum entanglement swapping network (see Fig. 1) is a scenario of three parties consisting of Alice, Bob and Charles, and two sources $\rho_{AB}$, $\rho_{BC}$ shared between them. All the three parties can perform measurements (or inputs) on their system, labeled $x$, $y$ and $z$ for Alice, Bob and Charles. After measuring the three parties, they obtain outcomes denoted $a$, $b$ and $c$, respectively. Here, $x,y,z,a,b,c\in\{0,1\}$. Bob’s measurement $y$ might correspond to a joint measurement on the two systems that he receives from each source. The correlations between the measurement outcomes of the three parties are described by the joint probability distribution $p(abc|xyz)$. In this scenario, the joint probability distribution $p(abc|xyz)$ is bilocal if it can be written as
$p(abc|xyz)=\int\int d\lambda d\mu q_{1}(\lambda)q_{2}(\mu)p(a|x,\lambda)p(b|y,\lambda,\mu)p(c|z,\mu),$
where $\lambda$ ($\mu$) characterizes the joint hidden variable of the system produced by the source $\rho_{AB}$ ($\rho_{BC}$) [70, 71].
Let $A_{x},B_{y},C_{z}$ denote the measurements of Alice, Bob, and Charles with inputs $x$, $y$, $z$, respectively. Note that the middle party Bob always performs joint measurement. Then, the bilocal inequality can be written as:
@@
where
@@
and $\langle A_{x}B_{y}C_{z}\rangle=\sum_{a,b,c=0}^{1}(-1)^{a+b+c}p(abc|xyz)$ [72].
In addition, one can choose different observables to construct different bilocal inequalities. For instance, the violation of the inequality
@@
can also be used to detect nonbilocal correlations, where $K=\langle A_{0}(B_{0}-B_{1})C_{0}\rangle$ and $L=\langle A_{1}(B_{0}+B_{1})C_{1}\rangle$ [73]. Of course, different inequalities exhibit varying capabilities in detecting nonlocality.
Mutually-commuting von Neumann algbera models. In QFT, the observables for Alice, Bob and Charles are associated with three mutually-commuting von Neumann algebras $\mathcal{M}_{A},\ \mathcal{M}_{B},\ \mathcal{M}_{C}$. In general, they are all required to be type III von Neumann algebra factors. Recall two von Neumann algebras satisfy Haag duality means that each algebra equals the commutant of the other. Haag duality encode quantum nonlocality intrinsically within their infinite-dimensional non-abelian structure through their type III factor property, which enforces that local algebras of spacelike separated regions are entangled with trivial relative commutants [74, 75]. Here, we do not restrict von Neumann algebras to III type factors, and generalize the Haag duality condition to mutual commutativity, proposing a von Neumann algebra model with mutual commutativity. Therefore, our model covers both scenarios of non-relativistic quantum mechanics and quantum field theory. The idea of this model is similar to that in Refs. [69, 68].
Definition 1. (Ternary mutually-commuting von Neumann algebra models) Let $\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C}$ be von Neumann subalgebras of $\mathcal{B}(H)$ over some Hilbert space $H$, and they are mutually commuting, i.e., $\mathcal{M}_{i}\subset\mathcal{M}_{j}^{\prime}$ with $i\neq j\in\{A,B,C\}$, where $\mathcal{M}_{j}^{\prime}$ is the communicator of $\mathcal{M}_{j}$. The generated von Neumann algebra
$\mathcal{M_{ABC}}=(\mathcal{M_{A}}\vee\mathcal{M_{B}}\vee\mathcal{M_{C}})^{\prime\prime}.$
We refer to the above model as the MCvNA model. When $\mathcal{M}_{ABC}=\mathcal{M}_{A}\otimes\mathcal{M}_{B}\otimes\mathcal{M}_{C}$, it is called the tensor product algebra model. Here, in order to correspond to the Hermitian property of quantum observables, these three algebras are always restricted to consist of self-adjoint elements.
We point out that the von Neumann algebra model defined above is intended for the study of algebraic structural properties; therefore, we have adopted the most general definition, which encompasses but does not correspond to models of quantum mechanics or quantum field theory.
We intend to use the above model of three mutually commuting von Neumann algebras to describe the entanglement swapping network in Fig. 1. We note that there is no correlation between parties $A$ and $C$ in the network. Mathematically, this independence can be described by the following formula. Let $\tau$ be a state in the dual space $\mathcal{M}_{ABC}^{*}$. In correspondence with the bilocal scenario, we always require that
@@
for any $a\in\mathcal{M}_{A},c\in\mathcal{M}_{C}$. We call it as the independent condition (*). This assumption will be used throughout this paper.

### v2

# Violation of Quantum Bilocal Inequalities on Mutually-Commuting von Neumann Algebra Models

@@ Abstract.
In contrast to the non-relativistic quantum mechanics, the violation of Bell inequalities in quantum field theory depends more on the structure of observable algebras (typically type III von Neumann algebras) rather than the choice of specific quantum states. Therefore, studying the violation of Bell inequalities based on the von Neumann algebraic framework can be applied to addressing questions in quantum field theory, while also providing insights into the structural properties of von Neumann algebras. In this paper, we employ three mutually-commuting von Neumann algebras to characterize quantum entanglement swapping networks, and establish Bell-like inequalities thereon, commonly referred to as bilocal inequalities. We investigate the algebraic structural conditions under which bilocal inequalities are satisfied or violated on the generated algebra of these three von Neumann algebras. Furthermore, the conditions for maximal violation of the inequalities can be utilized to infer the structural information of von Neumann algebras in reverse.
@@ 1. Introduction
In non-relativistic quantum mechanics, Bell nonlocality demonstrates that local measurements performed on one subsystem of a quantum state can instantaneously influence the measurement outcomes on another subsystem, regardless of the spatial separation between them [1, 2, 3]. Such nonlocal correlations can be detected through Bell inequalities, which serve as constraints that all local correlations must obey [4, 5, 6, 7, 8]. It has been demonstrated to offer quantum advantages in various device-independent quantum information tasks, including communication complexity [9], quantum key distribution [10, 11], randomness amplification [12, 13], and measurement-based quantum computation [14, 15]. However, it should be pointed out that relying solely on the non-relativistic quantum mechanics to discuss quantum information problems has two drawbacks [16, 17, 18]: 1) it fails to provide a universal framework for accurately describing phenomena in systems with infinite degrees of freedom; 2) it is not suitable for studying traditional quantum physics such as quantum field theory, which requires the language of type III von Neumann algebras. Research on quantum information problems on von Neumann algebras has received significant attention and yielded many meaningful results [19, 20, 21, 22, 23, 24, 25].
@@
In the 1980s early, mathematician Summers et al. introduce firstly the maximal violation of Bell inequalities and prove that its value is bounded by $2\sqrt${2} in the von Neumann algebra model, with equality attainable iff each subalgebra contains a copy of $\mathcal{M}_{2}({\mathbb{C}})$ [26]. This shows that Bell nonlocality is not merely a quantum peculiarity but a structural feature encoded in the classification of operator algebras, providing rigorous tools to quantify non-classical correlations in relativistic quantum systems [27]. Translating these bounds into the vacuum representation of algebraic quantum field theory, they show that tangent wedge algebras are always maximally correlated, whereas strictly spacelike-separated wedges decay exponentially with mass-governed distance [28, 29, 30, 31, 32]. These works reveal a novel algebraic invariant, termed the Bell correlation invariant, which distinguishes infinitely many isomorphism classes of pairs of mutually commuting von Neumann algebras and links the maximal violation to the occurrence of the hyperfinite type $\rm II_{1}$ factor [27]. This is a pioneering work to make Bell nonlocality in quantum field theory (QFT) serve as a crucial bridge connecting quantum information science with fundamental physics [27, 33]. It provides a rigorous framework for reconciling quantum entanglement with relativistic causality, resolves conceptual challenges such as impossible measurements, and reveals how fundamental symmetries like parity violation affect quantum correlations [34, 35, 36, 37, 38].
@@
In contrast to entanglement originating from an individual source, quantum networks comprise numerous small-scale entangled states. Owing to the independence among distinct sources, the correlations emerging from quantum networks exhibit non-convex characteristics that transcend the polytopes associated with single-source entanglement [39, 40, 41, 42, 43, 44]. To date, Bell-like inequalities in the non-relativistic quantum mechanics have been devised to certify nonlocal correlations across diverse network architectures, such as entanglement-swapping networks [39, 45], chain configurations [46, 47], star topologies [48, 49], polygon structures [50, 51, 52], tree-shaped networks [53, 54, 55], arbitrary acyclic networks [41, 42, 56], and arbitrary $k$-independent networks [57]. Alternative research directions examine the stronger forms of network nonlocality that surpass hybrid implementations involving classical variables and post-quantum resources [58, 59]. Nevertheless, limited progress has been made concerning the discrimination of correlations produced by different networks and the subsequent identification of underlying quantum network topologies [60]. Recently, the notion of bi-locality in an entanglement swapping network based on the mutually-commuting von Neumann algebra model has already been introduced by Ligthart et al. [61], and Xu has addressed the inclusion problem between the tensor product algebra model and the mutually-commuting von Neumann algebra model in this setting [62]. However, Bell-like inequalities for such von Neumann algebraic models have not yet been established. In this paper, we aim to establish bi-local inequalities within the mutually-commuting von Neumann algebra model and investigate how the degree of their violation is related to the structural properties of the algebras.
2. Mutually-commuting von Neumann algebra models, Tsirelson problem, and Connes’ embedding conjecture
In QFT, the algebras assigned to touching spacetime regions are of type III von Neumann ones, so there are no pure normal states and no tensor-product structures [63, 64, 65]. This differs from the non-relativistic quantum-mechanical setup, which is usually linked to type I von Neumann algebras and relies on the algebraic tensor product as its mathematical framework [66, 67, 68]. These are two distinct models, referred to respectively as the tensor product algebra model and the mutually-commuting von Neumann algebra model.
In the tensor-product algebra model, a bipartite composite quantum system is described by the tensor product of two Hilbert spaces $\mathcal{H}_{\mathcal{A}}$ and $\mathcal{H}_{\mathcal{B}}$, denoted as $\mathcal{H}=\mathcal{H}_{\mathcal{A}}\otimes\mathcal{H}_{\mathcal{B}}.$ Let $\mathcal{B}(\mathcal{H})$ denote the algebra of all bounded linear operators on $\mathcal{H}$, and $\mathcal{T}(\mathcal{H})$ the space of trace-class operators. Then there is a duality relation $\mathcal{B}(\mathcal{H})=\mathcal{T}(\mathcal{H})^{*}$. A quantum state $\rho$ is a positive semidefinite operator with trace 1, and the set of states is denoted by $\mathcal{S}(\mathcal{H})=\{\rho\in\mathcal{T}(\mathcal{H})\mid\rho\geq 0,\ \mathrm{Tr}(\rho)=1\}$. The algebra of observables is the Jordan algebra of self-adjoint operators, denoted by $\mathcal{B}_{s}(\mathcal{H})=\{A\in\mathcal{B}(\mathcal{H})\mid A=A^{*}\}.$ In the mutually-commuting von Neumann algebra model, the algebra of observables of a quantum system is described by a von Neumann algebra $\mathcal{M}$, with $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{B}}$ being two mutually commuting von Neumann subalgebras of $\mathcal{M}$ such that $(\mathcal{M}_{\mathcal{A}}\vee\mathcal{M}_{\mathcal{B}})^{\prime\prime}=\mathcal{M}$. Here, $\mathcal{M}^{\prime\prime}$ denotes the double commutant of $\mathcal{M}$.
However, in more general cases, a natural question arises: are the two models isomorphic? In other words, does $\mathcal{M}$ admit an isomorphism to $\mathcal{M}_{\mathcal{A}}\otimes\mathcal{M}_{\mathcal{B}}$? This question is intimately related to the famous Tsirelson problem, which is equivalent to the Connes embedding conjecture in von Neumann algebra theory [69, 70, 71]. The Tsirelson problem asks whether the sets of joint probability distributions of measurement outcomes are the same in the two models. In 2020, Ji and collaborators [71] designed a quantum interactive protocol based on the classical halting problem in computability theory, and proved that the answer to the Tsirelson problem is negative. This means that the tensor-product algebra model and the mutually-commuting von Neumann algebra model are not equivalent. Therefore, the mutually-commuting von Neumann algebra model provides a more general framework, and many novel quantum phenomena fundamentally different from those in the tensor-product model have been discovered in this setting [71, 72, 73, 74].
@@ 3. Ternary mutually-commuting von Neumann algebra models and entanglement swapping networks
Quantum bilocal scenario. In non-relativistic quantum mechanics, the quantum entanglement swapping network (see Fig. 1) is a scenario of three parties consisting of Alice, Bob and Charles, and two sources $\rho_{AB}$, $\rho_{BC}$ shared between them. The inputs and outputs of the measurements performed by the three parties are denoted as $x,y,z$ and $a,b,c$, respectively, where $x,y,z,a,b,c\in\{0,1\}$. Assume that each party performs binary-input and binary-output measurements, with the observables for Alice, Bob, and Charles denoted as $A_{x}$, $B_{y}$ and $C_{z}$, respectively. Here it is required that the spectra of operators $A_{x},B_{y},C_{Z}$ are all {-1,1}, implying that $I\leq A_{x},B_{y},C_{z}\leq I$. The correlations between the measurement outcomes of the three parties are described by the joint probability distribution $p(abc|xyz)$. In this scenario, $p(abc|xyz)$ is said to be bilocal if it can be written as
$p(abc|xyz)=\int\int d\lambda d\mu p_{1}(\lambda)p_{2}(\mu)p(a|x,\lambda)p(b|y,\lambda,\mu)p(c|z,\mu),$
where $\lambda$ and $\mu$ characterize the hidden variables of the systems produced by the sources $\rho_{AB}$ and $\rho_{BC}$, respectively [75, 76]. Otherwise, it is called non-bilocal.
In order to detect non-bilocal correlations generated by the network, it is often necessary to find suitable measurements that violate the following bilocal inequality
@@
whose maximum quantum violation is $2\sqrt{2}$ and is attainable. Here
@@
as introduced in Ref. [77]:
$\langle A_{x}B_{y}C_{z}\rangle=\sum_{a,b,c=0}^{1}(-1)^{a+b+c}{\rm tr}((A_{a|x}B_{b|y}C_{x|z})\rho_{AB}\otimes\rho_{BC})=\sum_{a,b,c=0}^{1}(-1)^{a+b+c}p(abc|xyz).$
Here $A_{x}=\sum_{a}(-1)^{a}A_{a|x}$, $B_{y}=\sum_{b}(-1)^{b}B_{b|y}$ and $C_{z}=\sum_{z}(-1)^{z}C_{c|z}$, where $A_{a|x},\ B_{b|y}$, and $C_{c|z}$ are the positive operator-valued measurement (POVMs) performed by Alice, Bob and Charles, respectively.
Furthermore, it follows from Ref. [57] that the violation of the inequality
@@
can also be used to detect nonbilocal correlations, where $K=\langle A_{0}(B_{0}-B_{1})C_{0}\rangle$ and $L=\langle A_{1}(B_{0}+B_{1})C_{1}\rangle$ [78]. Notably, different inequalities exhibit varying capabilities in detecting nonlocality.
Mutually-commuting von Neumann algebra models. In QFT, the observables for Alice, Bob, and Charles are associated with three mutually-commuting von Neumann algebras $\mathcal{M}_{\mathcal{A}}$, $\mathcal{M}_{\mathcal{B}}$, $\mathcal{M}_{\mathcal{C}}$. Therefore, our model encompasses both the non-relativistic quantum mechanics scenario and the quantum field theory scenario. The idea of this model is similar to that in Refs. [61, 62].
Definition 1. (Ternary Mutually-commuting von Neumann Algebra Models of Tripartite Quantum Systems) Let $\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}}$ be von Neumann subalgebras of $\mathcal{B}(\mathcal{H})$ over some Hilbert space $\mathcal{H}$, which are mutually commuting, i.e., $\mathcal{M}_{i}\subset\mathcal{M}_{j}^{\prime}$ with $i\neq j\in\{\mathcal{A},\mathcal{B},\mathcal{C}\}$, where $\mathcal{M}_{j}^{\prime}$ is the communtant of $\mathcal{M}_{j}$. The generated von Neumann algebra
$\mathcal{M}_{\mathcal{ABC}}=(\mathcal{M}_{\mathcal{A}}\vee\mathcal{M}_{\mathcal{B}}\vee\mathcal{M}_{\mathcal{C}})^{\prime\prime}.$
We refer to the above model as the TMCvNA model. When $\mathcal{M}_{\mathcal{ABC}}\simeq\mathcal{M}_{\mathcal{A}}\otimes\mathcal{M}_{\mathcal{B}}\otimes\mathcal{M}_{\mathcal{C}}$, it is called the tensor product algebra model. In this paper, for any $A\in\mathcal{M}_{\mathcal{A}}$, $B\in\mathcal{M}_{\mathcal{B}}$, $C\in\mathcal{M}_{\mathcal{C}}$, we always assume that they are Hermitian.
We intend to use the above model of ternary mutually commuting von Neumann algebras to describe the entanglement swapping network in Fig. 1. We note that there is no correlation between the parties $\mathcal{A}$ and $\mathcal{C}$ in the network. Mathematically, this independence can be described by the following formula. So the network state of the entanglement swapping network $\tau$ should be a state in the dual space $\mathcal{M}_{\mathcal{ABC}}^{*}$, satisfying
@@
for any $A\in\mathcal{M}_{\mathcal{A}},C\in\mathcal{M}_{\mathcal{C}}$. We call it the independent condition (*). This assumption will be used throughout this paper.
Definition 2. A ternary mutually-commuting von Neumann algebra model of entanglement swapping networks is the ternary mutually-commuting von Neumann algebra model of tripartite quantum systems with the independence condition (*).